OpenStudy (anonymous):

Find the equation of a line with the given slope and containing the point m=0, (7, -1)

OpenStudy (katrinakaif):

Put it into: (y-y1) = m (x- x1) as (7,-1) is (x1,y1)

OpenStudy (unklerhaukus):

m = o !

OpenStudy (anonymous):

y=-1

OpenStudy (amistre64):

y = mx -mPx + Py y = 0x -0(7) +-1 y = -1